OpenStudy (anonymous):

Solve the system of equations for x or y, whichever one you do not eliminate. 2x+4y=10 2x-4y=10 a.x=2 b.x=10 c.x=5 d.y=5

sam (.sam.):

2x+4y=10 2x-4y=10 ------------ 8y=0 y=0 2x-4y=10 2x-4(0)=10 2x=10 x=5

OpenStudy (anonymous):

Thanks friend =)